OpenStudy (anonymous):

A plane travels from Orlando to Denver and back again. On the five-hour trip from Orlando to Denver, the plane has a tailwind of 40 miles per hour. On the return trip from Denver to Orlando, the plane faces a headwind of 40 miles per hour. This trip takes six hours. What is the speed of the airplane in still air?

OpenStudy (mathstudent55):

Let the speed of the plane in still air = x When there is a tailwind, the speed is x + 40 When tere is a headwind, the speed is x - 40 speed = distance/time, or s = d/t distance = speed * time, or d = st Orlando to Denver d = st = (x + 40)5 Denver to Orlando d = st = (x - 40)6 The distance is the same both ways, so (x + 40)5 = (x - 40)6 Solve the equation for x.